OpenStudy (anonymous):

Simplify the expression. 3^7/3^5

OpenStudy (mathstudent55):

Here's a rule you need to memorize: \( \dfrac{a^m}{a^n} = a^{m - n} \) Here's an example of the rule being used: \( \dfrac{5^{12}}{5^5} = 5^{12 - 5} = 5^7 \) Now apply the rule to your problem.

OpenStudy (anonymous):

so 3^2?

OpenStudy (anonymous):

If so that isn't an answer choice

OpenStudy (mathstudent55):

That is correct; you need to do the next step. What does 3^2 mean?

OpenStudy (anonymous):

three squared

OpenStudy (mathstudent55):

Right. What number is 3 squared?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

ha, so 9 is the answer :D

OpenStudy (anonymous):

Would you mind helping me with one more ?

OpenStudy (mathstudent55):

correct

OpenStudy (mathstudent55):

I will, but pls start a new post for a new problem.

OpenStudy (anonymous):

Okay will do